A volume of 20 litres of hydrogen chloride was completely decomposed to its constituent element
(hydrogen and chloride). what are the volumes of this constituent elements
1
Expert's answer
2020-04-02T08:56:38-0400

Reaction:

2HCl = H2 + Cl2

from the equation we can find out:

n(H2) = n(Cl2) = "{\frac {n(HCl)} {2}}"

In same conditions molar volumes of gases are same too, and if n*Vm = V then

"{\frac {V(H_2)} {V_m}}= {\frac {V(Cl_2)} {V_m}}= {\frac {V(HCl)} {2V_m}}"

"V(H_2)=V(Cl_2)= {\frac {V(HCl)} {2}}={\frac {20} {2}}=10" litres

V(H2) = 10 litres

V(Cl2) = 10 litres